# Threshold Graphs are Globally Synchronizing

### v1

@@ 4.3. Force and stability
Figure 7. $\boldsymbol{v}_{i}$ is pinned down at zero angle, which determines the region where $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\boldsymbol{v}_{j}$ should be (the right-end unrestricted region). The dashed vectors represent the phasors of nodes connected to ii.
Proof of Lemma 4.4.
We begin by noting that the $i$-th diagonal entry of the Hessian matrix $\mathbf{H}$ at $(\theta_{1},\cdots,\theta_{n})$ can be rewritten as
$\mathbf{H}_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
where the second equality uses the equilibrium condition (4.1), and the last equality is deduced from re-writing the unit vectors in condition $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}$ into the form using complex numbers. If there exists $\mu_{i}<0$, then equivalently $\mathbf{H}_{ii}<0$, which contradicts the positive definiteness of $\mathbf{H}$, see [Bha07] [HJ85]. Hence, $\boldsymbol{\theta}$ cannot be a second-order critical point. Consequently, it is neither a local minimum of (3) nor a stable equilibrium of (2). Therefore, $\mathbf{H}_{ii}=\mu_{i}>0$ holds for all $i\in[n]$, where $n$ is the size of $G$. ∎
To summarize, the lemma above shows that at any second-order critical point $\boldsymbol{\theta}$, the phasor of each node must be positively aligned with the sum of the phasors of its neighbors (whenever the latter is nonzero). That is,
$\boldsymbol{v}_{i}\;\Uparrow\;\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}=\boldsymbol{0},$
for all $i\in[n]$.
Notation. Here, We write $a\Uparrow b$, if both are nonzero and point in the same direction (i.e., the angle between them is zero). In the remainder of this manuscript, we will use this notation, as we find it more intuitive.
Remark 4.6 (This is only a necessary condition).
Conversely, at a critical point (equilibrium), all $\mu_{i}>0$ for all $i\in V$ is not sufficient to guarantee that $\boldsymbol{\theta}$ is a second-order critical point. As a counter-example, the equilibrium $\boldsymbol{\theta}^{*}=(0,0,0,0,\pi,\pi,\pi,\pi)$ of the Kuramoto model on the graph in Figure 8 is a SOSP.
12345678
Figure 8. At the configuration $(0,0,0,0,\pi,\pi,\pi,\pi)$, each node experiences a force from its neighbors that is aligned with its own direction. However, this configuration is not stable, as can be verified by examining the eigenvalues of the corresponding Hessian matrix.
Indeed, the configuration $(0,0,0,0,\pi,\pi,\pi,\pi)$ is not a local minimum of the energy function. When nodes $1,2,3,4$ are jointly perturbed slightly away from angle 0 in the same direction, they do not return to the original configuration. Instead, they continue to move toward the $\pi$-oriented nodes $5,6,7,8$, indicating a decrease in energy and confirming that the configuration is unstable. This can be rigorously proven by quadratic form $\boldsymbol{x}^{T}\boldsymbol{H}\boldsymbol{x}$, we omit it here.
4.4. Notations
The below definitions are nodewise concept since they are always made with respect to a node in certain state $\boldsymbol{\theta}$. For example, when we say that node $i$ attains equilibrium at $\boldsymbol{\theta}$, we mean that
$\mu_{i}\boldsymbol{v}_{i}\;=\;\sum_{j\in N(i)}\boldsymbol{v}_{j},\qquad\mu_{i}\in\mathbb{R}.$
Similarly, when we say that node $i$ attains a stable equilibrium at $\theta$, we mean the same balance equation (8) holds with $\mu_{i}\geq 0$. In the proof of global synchrony for threshold graphs, we frequently use this necessary condition. To make the reasoning more intuitive for the reader, we assign different colors to nodes representing different cases.
Condition on $i$ Equation Illustration
Equilibrium $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}\in\mathbb{R}$ $i$
Stable equilibrium $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}\geq 0$ $i$
Asymptotically stable (positively forced) $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}>0$ $i$
Zero-forced $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}=0$ $i$
Negatively forced $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}<0$ $i$
Table 1. Stability classification of node $i$ at equilibrium $\theta$. The sign of $\mu_{i}$ determines the type of stability, visualized by node color.
It is important to note that these definitions do not require the state $\boldsymbol{\theta}$ itself to be an equilibrium or stable equilibrium of the system. Indeed, at a given state $\boldsymbol{\theta}$ it is possible that some nodes satisfy the condition that their own phasor is parallel to the aggregated phasor formed by their neighbors. Such nodes are in equilibrium at this particular configuration $\boldsymbol{\theta}$; however, this does not imply that the configuration itself is an equilibrium state, since other nodes may not satisfy the same condition. Nevertheless, if $\boldsymbol{\theta}$ is indeed such an equilibrium, then the above conditions are of course implied, according to Lemmas 4.4 and 4.1.
5. Geometric twins and their phasor geometry
In this section, we introduce the concept of geometric twins (g-twins for abbreviation), that takes into account the unit vectors $\boldsymbol{v}_{i}=(\cos\theta_{i},\sin\theta_{i})$ associated with each node $i$ in a Kuramoto state $\boldsymbol{\theta}$. Two nodes are said to be g-twins if they satisfy certain geometric relationships at this configuration. It is worth noting that, due to the conceptual similarity, we borrow the term twins from graph theory. However, we precede it with the word geometric to emphasize that our definition depends on the positions of the node phasors in the specific configuration. This concept captures a form of geometric symmetry at $\boldsymbol{\theta}$ that will be crucial in our proof for global synchrony of threshold graphs.
5.1. Closed geometric twins
Definition 5.1 (Closed g-twins).
Let $G=(V,E)$ be a graph and $\boldsymbol{\theta}$ a state of the Kuramoto model on $G$. Two nodes $i,j\in V$ are called closed g-twins at state $\boldsymbol{\theta}$ if there exists a two-dimensional vector $\boldsymbol{q}$ such that
$\bigl(\boldsymbol{v}_{i}\parallel\boldsymbol{v}_{j}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{j}+\boldsymbol{q}=\boldsymbol{0}\bigr)\\ \quad\text{and}\quad\bigl(\boldsymbol{v}_{j}\parallel\boldsymbol{v}_{i}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{i}+\boldsymbol{q}=\boldsymbol{0}\bigr)$
or equivalently
$\boldsymbol{v}_{j}+\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}\quad\text{and}\quad\boldsymbol{v}_{i}+\boldsymbol{q}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\in\mathbb{R}$
where $\boldsymbol{v}_{k}=(\cos\theta_{k},\sin\theta_{k})$ is the unit vector associated with node $k$.
If moreover both strengths are nonnegative, meaning that each vector $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ points in the same direction as the sum of the other and the common vector $\boldsymbol{q}$ (or that summation is the zero vector), then we call them stable closed g-twins.
Theorem 5.2 (Phasor geometry of closed g-twins).
Consider a state $\boldsymbol{\theta}$ of the Kuramoto model on graph $G=(V,E)$, the phasor positions of closed g-twins $a,b\in V$ with strength $\mu_{a},\mu_{b}$ with respect to the common vector falls into one of the following three cases:
@@ 5.1. Closed geometric twins
    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$ and $\mu_{a}=\mu_{b}=-1$ and $\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=0$.
$\mu_{a}$$\mu_{b}$
Figure 9. Feasible region of strengths $(\mu_{a},\mu_{b})$ for closed g-twins. It lies on the lines $\mu_{a}=\mu_{b}$ (black, with $a$ and $b$ synchronized) and $\mu_{a}+\mu_{b}=-2$ (red, with $a$ and $b$ in antipodal position). At their intersection (1,1)(-1,-1), marked by a hollow dot, there is no additional constraint on the phasors of $a$ and $b$.
@@
According to the definition of closed g-twins
$\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}\quad\text{and}\quad\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$
we have
@@
Definition 5.3 (Stable closed g-twins).
Under the setting of Definition 5.1, we say that nodes $i$ and $j$ are stable closed g-twins if furthermore
$\bigl(\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}_{j}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{j}+\boldsymbol{q}=\boldsymbol{0}\bigr)\\ \quad\text{and}\quad\bigl(\boldsymbol{v}_{j}\Uparrow\boldsymbol{v}_{i}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{i}+\boldsymbol{q}=\boldsymbol{0}\bigr)$
or equivalently
$\boldsymbol{v}_{j}+\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}\quad\text{and}\quad\boldsymbol{v}_{i}+\boldsymbol{q}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\geq 0.$
Similarly, one can define open g-twins and stable open g-twins at state $\boldsymbol{\theta}$, though this concept will not be used in our proofs in this chapter. It may nevertheless be of independent interest in understanding the geometry of non-adjacent nodes with certain geometric symmetries at equilibria.
5.2. Open geometric twins
Definition 5.4 (Open g-twins).
Let $G=(V,E)$ be a graph and $\boldsymbol{\theta}$ a state of the Kuramoto model on $G$. Two nodes $i,j\in V$ are called open g-twins at state $\boldsymbol{\theta}$ if there exists a two-dimensional vector $\boldsymbol{q}$ such that
$\left(\boldsymbol{v}_{i}\parallel\boldsymbol{q}\quad\text{and}\quad\boldsymbol{v}_{j}\parallel\boldsymbol{q}\right)\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
or equivalently,
$\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\in\mathbb{R}.$
Here $\boldsymbol{v}_{k}=(\cos\theta_{k},\sin\theta_{k})$ is the unit vector associated with node $k$.
Definition 5.5 (Stable open g-twins).
Under the setting of Definition 5.4, we say that nodes $i$ and $j$ form geometric stable open twins if furthermore
$\boldsymbol{v}_{i}\Uparrow\boldsymbol{q}\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
and
$\boldsymbol{v}_{j}\Uparrow\boldsymbol{q}\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
or equivalently,
$\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\geq 0.$
Remark 5.6.
For brevity, whenever the conditions of Definition 5.4 are satisfied, we shall simply say that $i$ and $j$ form open g-twins with strengths $(\mu_{i},\mu_{j})$, with respect to the common vector $\boldsymbol{q}$. The terminology for closed g-twins, stable g-twins, and stable closed g-twins is used analogously.
Under the special geometric symmetries described above, the phasors of g-twins exhibit a particular symmetric configuration on the unit circle. This in turn allows us to obtain a precise characterization of the relative positions of the phases $\theta_{i}$ of nodes with such hidden geometric symmetry, both at equilibrium and at stable equilibrium of the Kuramoto model. We refer to the notation on node at (stable) equilibrium in Table 1 for clarity.
Theorem 5.7 (Phasor geometry of open g-twins).
Consider a state $\boldsymbol{\theta}$ of the Kuramoto model on graph $G=(V,E)$, the phasor positions of open g-twins $a,b\in V$ with strength $\mu_{a},\mu_{b}$ with respect to the common vector $\boldsymbol{q}$ falls into one of the following three cases:
  1. (1)
    $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$ and $\mu_{a}=\mu_{b}\neq 0$;
  2. (2)
    $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$ and $\mu_{a}=-\mu_{b}\neq 0$;
  3. (3)
    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$ and $\mu_{a}=\mu_{b}=0$.
@@ 5.2. Open geometric twins
Figure 10. Feasible region of strengths $(\mu_{a},\mu_{b})$ for open g-twins. It lies on the lines $\mu_{a}=\mu_{b}$ (black, corresponding to $a$ and $b$ synchronized) and $\mu_{a}=-\mu_{b}$ (red, corresponding to $a$ and $b$ antipodal). At their intersection $(0,0)$, marked by a hollow dot, there is no additional constraint on the phasors of $a$ and $b$.
@@
Since
$\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}=\mu_{b}\boldsymbol{v}_{b},$ (14)
the following cases arise.
Case 1. $\mu_{a}=\mu_{b}=0$, then we have $\boldsymbol{q}=0$.
Case 2. If $\mu_{a},\mu_{b}\neq 0$, then, from 14 we have
$\boldsymbol{v}_{a}=\frac{1}{\mu_{a}}\boldsymbol{q}=\frac{\mu_{b}}{\mu_{a}}\boldsymbol{v}_{b}.$
Since $\|\boldsymbol{v}_{a}\|=\|\boldsymbol{v}_{b}\|=1$, it follows that $|\mu_{a}|=|\mu_{b}|$. We distinguish two subcases. If $\mu_{a}=\mu_{b}$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. If $\mu_{a}=-\mu_{b}$, then $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$, and one of $\mu_{a},\mu_{b}$ is negative. ∎
6. Formation of geometric twins
The motivation for introducing g-twins in the previous section comes from the proof of global synchronization on threshold graphs. In that proof, it turns out that synchronization does not rely solely on structural symmetry (i.e., forming structural twins), but can also arise from a more general geometric symmetry, where the corresponding phasors occupy special symmetric positions on the unit circle.
In this section, we first mention that structural twins at equilibrium constitute the simplest case of forming g-twins. We then move on to more general situations where nodes are not structural twins, yet under both specific geometric conditions (on the phasors) and accompanying structural conditions, they still form g-twins. Both cases will play a crucial role in the proof of global synchronization for threshold graphs.
6.1. Structural twins at equilibrium
Definition 6.1 (Structural Twins).
Let $G=(V,E)$ be a graph. Two nodes $i,j\in V$ are called structural closed twins if
@@ 6.1. Structural twins at equilibrium
and structural open twins if
$N(i)=N(j),$ (16)
(a) Closed twins ($N[i]=N[j]$).
(b) Open twins ($N(u)=N(v)$).
Figure 11. Illustration of structural twins.
Remark 6.2.
In the discussion above we used the term structural to emphasize that whether two nodes are twins depends purely on the graph structure. In graph theory they are simply called true or adjacent twins, and false or non-adjacent twins.
Clearly, if $\boldsymbol{\theta}$ is an equilibrium point of the Kuramoto model with respect to graph $G$, and two nodes $a$ and $b$ satisfy condition 15, then they must form a pair of closed g-twins. This is because, according to Lemma 4.1, the nodes $a$ and $b$ are structurally symmetric (satisfying (15)) 8, that is
[…]
$\boldsymbol{q}=\sum_{j\in N(a)\setminus\{b\}}\boldsymbol{v}_{j}=\sum_{j\in N(b)\setminus\{a\}}\boldsymbol{v}_{j}$
such that $\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$ for some $\mu_{a},\mu_{b}\in\mathbb{R}$, according to notations in 4.4. The relative positions of nodes $a$ and $b$ on the unit circle, together with their nodewise stability in this case, are illustrated in Figure 14.
There, colors indicate node-wise stability: Blue indicates $\mu>0$ (stable), red for $\mu<0$ (unstable), black for $\mu=0$.
Figure 12. Structural closed twins attaining nodewise equilibrium.
Remark 6.3.
Here the case (d) is detailed in the following figure. One can clearly see that
$\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=\boldsymbol{0},$
where $N$ denotes the set of common neighbors of $a$ and $b$, excluding each other.
Figure 13. Geometric illustration of case (d) for false twins at equilibrium. The red arrow represents the vector sum $\boldsymbol{q}$, where $N$ is the common neighborhood of nodes $a$ and $b$, excluding each other. In this case, $\mu_{a}=\mu_{b}=-1$, and the three vectors $\boldsymbol{v}_{a}$, $\boldsymbol{v}_{b}$, and $\boldsymbol{q}$ form a triangle summing to zero.
Similarly, structural open twins $a,b$ at equilibrium in $\boldsymbol{\theta}$ are also geometric open twins, since there exists
$\boldsymbol{q}=\sum_{j\in N(a)}\boldsymbol{v}_{j}=\sum_{j\in N(b)}\boldsymbol{v}_{j}$
such that $\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}$ and $\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$ for some $\mu_{a},\mu_{b}\in\mathbb{R}$, according to notations in 4.4. The relative positions of nodes $a$ and $b$ on the unit circle, together with their nodewise stability in this case, are illustrated in Figure 14. The same as in the above mentioned closed twins case, that colors indicate their stability.
Figure 14. Structural open twins attaining nodewise equilibrium.
Our strategy for proving that a given graph admits global synchronization is to characterize the positions of all nodes at stable equilibrium (in fact, more general, at SOSP of $E(\boldsymbol{\theta})$). Furthermore, the lemma on structural twins at equilibrium shows that whenever two structural twins both attain stable equilibrium, they must synchronize if they are closed twins. It is worth noting, however, that in the case of open twins, synchronization is not guaranteed: it may happen that both strengths $\mu_{i}$ and $\mu_{j}$ vanish, in which case the phasors of $i$ and $j$ remain unconstrained.
At this point, we are already able to characterize certain new classes of globally synchronizing graphs, according to results in this section.
@@
Complete split graphs, windmill graphs, complete bipartite graphs are globally synchronizing.
(a) Complete split graphs.
(b) Windmill graphs $W_{5,6}$.
(c) Complete bipartite.

### v2

@@ 4.3. Angle and Stability
Figure 7. With $\boldsymbol{v}_{i}$ fixed at angle 0, the endpoint of the aggregated phasor $\sum_{j}A_{ij}\boldsymbol{v}_{j}$ is constrained to the right-end sliding track. The rods may freely adjust their orientations, but the endpoint of the last rod must lie in the feasible sliding interval. The solid vectors show the unit-length phasors of the neighbors of node ii.
The geometric conditions of second-order stationary points presented here are only the most basic one. Further necessary geometric conditions will be established in subsequent sections, and these will be essential for showing that every connected threshold graph is globally synchronizing.
5. Local Synchronization Primitives
5.1. A Fundamental Geometric Fact
We begin with a geometric lemma that characterizes all possible relative positions of two unit vectors satisfying a pair of linear relations. This fact will serve as the local synchronization primitive used in the induction argument in Section 6.
Lemma 5.1.
Let $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}\subset\mathbb{R}^{2}$. Suppose there exist a vector $\boldsymbol{q}\in\mathbb{R}^{2}$ and scalars $\mu_{a},\mu_{b}\in\mathbb{R}$ such that
$\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}\qquad\text{and}\qquad\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}.$ (12)
Then the positions of $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$ fall into one of the following three cases:
@@ 5.1. A Fundamental Geometric Fact
    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$, $\mu_{a}=\mu_{b}=-1$, and $\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=0$.
@@
According to (12), we have
@@
Figure 8. The feasible region for $(\mu_{a},\mu_{b})$ consists of: (1) the black line $\mu_{a}=\mu_{b}$, excluding the intersection point, corresponding to $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) the red line $\mu_{a}+\mu_{b}=-2$, excluding the intersection point, corresponding to the antipodal case $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$; and (3) their intersection point (1,1)(-1,-1), marked by a hollow dot, at which no additional constraint is imposed on the vectors $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$.
Corollary 5.2.
Under the assumptions of Lemma 5.1, if in addition $\mu_{a},\mu_{b}\geq 0$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$.
@@
With $\mu_{a},\mu_{b}\geq 0$, Lemma 5.1 leaves only case (1), hence $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
Definition 5.3 (Geometric twins).
At state $\boldsymbol{\theta}$ of Kuramoto model on $G$, nodes $i$ and $j$ are called geometric twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the compatibility relations of Lemma 5.1.
5.2. Phasor Geometry of Closed Twins
Definition 5.4.
Let $G=(V,E)$ be a graph. Two nodes $i,j\in V$ are called closed twins if
@@ 5.2. Phasor Geometry of Closed Twins
where $N[i]$ is defined as $N(i)\cup\{i\}$.
Figure 9. The two black nodes form a pair of closed twins.
Open twins, defined by $N(i)=N(j)$, will not be needed in this work.
Corollary 5.5.
Let $\boldsymbol{\theta}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$. If nodes $a$ and $b$ form a pair of closed twins, then $\theta_{a}=\theta_{b}$.
Proof.
From Lemma 4.4, since $\boldsymbol{\theta}$ is a second-order stationary point, there exist $\mu_{a},\mu_{b}\geq 0$ such that
$\begin{cases}\displaystyle\sum_{j\in N(a)}\boldsymbol{v}_{j}=\mu_{a}\boldsymbol{v}_{a},\\[4.0pt] \displaystyle\sum_{j\in N(b)}\boldsymbol{v}_{j}=\mu_{b}\boldsymbol{v}_{b}.\end{cases}$
Because $N[a]=N[b]$, we have $N(a)\setminus\{b\}=N(b)\setminus\{a\},$ and we denote this common set by $M$. Thus,
$\begin{cases}\displaystyle\sum_{j\in M}\boldsymbol{v}_{j}+\boldsymbol{v}_{b}=\mu_{a}\boldsymbol{v}_{a},\\[4.0pt] \displaystyle\sum_{j\in M}\boldsymbol{v}_{j}+\boldsymbol{v}_{a}=\mu_{b}\boldsymbol{v}_{b}.\end{cases}$
By Corollary 5.2 and the fact that $\mu_{a},\mu_{b}\geq 0$, we conclude that $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$, and therefore $\theta_{a}=\theta_{b}$. ∎
Clearly, this is a case where geometric twins defined in (5.3) formed. Moreover, Corollary 5.5 implies immediately that windmill graphs are globally synchronizing.
@@
Every windmill graph is globally synchronizing.
Proof.
Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$, and let $G$ be a windmill graph. Within each clique of $G$, all non-central nodes have identical closed neighborhoods and therefore form closed twins. By Corollary 5.5, all nodes in each clique must share the same phase. Contracting each clique into a single node yields a star graph, which is globally synchronizing. Hence all nodes of $G$ have the same phase. ∎
Figure 10. Windmill graphs $W_{5,6}$.